Calculate the total amount of energy needed to change 2.57 g of water from an original temp of -73.2 C to a final temperature of 221 C.
1
Expert's answer
2020-02-28T12:12:06-0500

Ex=Cxmx"\Delta"T

E = E-73.2/0 + Emelt + E0/100 + Evap + E100/221 = 2.57 × 18.103 × 73.2 + 2.57 × 335.105 + 2.57 × 4.220 × 100 + 2260 × 2.57 + 2.57 × 2.000 × 221 = 12259 J = 12.26 kJ